OpenStudy (anonymous):

what is the solution set of x² + 8x + 12 = 0?

jimthompson5910 (jim_thompson5910):

Hint: x^2 + 8x + 12 factors to (x+6)(x+2)

OpenStudy (anonymous):

so is it (2,6)

jimthompson5910 (jim_thompson5910):

close, but not quite

OpenStudy (anonymous):

(-2,-6)?

jimthompson5910 (jim_thompson5910):

(x+6)(x+2) = 0 x+6=0 or x+2=0 x = ?? or x = ??

jimthompson5910 (jim_thompson5910):

yep, x = -2 or x = -6

OpenStudy (anonymous):

thank you so much !

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

what about.... equivalent form of the equation t²+5t-24=0?

jimthompson5910 (jim_thompson5910):

factor the left side

OpenStudy (anonymous):

alright then?...

jimthompson5910 (jim_thompson5910):

what do you get when you factor

OpenStudy (anonymous):

i have no idea honestly lol

jimthompson5910 (jim_thompson5910):

find two numbers that a) multiply to -24 AND b) add to 5

OpenStudy (anonymous):

ummm idk :(

jimthompson5910 (jim_thompson5910):

list out the factors of -24 and the possible ways to multiply to -24

jimthompson5910 (jim_thompson5910):

then add each pair of factors

OpenStudy (anonymous):

12 and 2...6 and 4...8 and 3

jimthompson5910 (jim_thompson5910):

since -24 is negative, one of the factors must be negative (the other must be positive)

OpenStudy (anonymous):

im so lost can you just tell me what it would look like ?...

OpenStudy (anonymous):

(t + 8)(t - 3) = 0 (t - 8)(t + 3) = 0 (t - 8)(t - 3) = 0

OpenStudy (anonymous):

those are the choices ?

jimthompson5910 (jim_thompson5910):

8 and -3 both a) multiply to -24 AND b) add to 5